The intensity of light L, measured in foot-candles, varies inversely as the square of the distance from the light d. if L=4/5 foot-candles when the distance is 20 feet. Find the constant of variation and then find the intensity when the distance is 16 feet.
K=___________
answer________________

Answer by mukhopadhyay(490)

L = k/d^2;
=> k = Ld^2
Using the values provided in the question,
k = (4/5)*(20^2) = 400*4/5 = 320
L = 320/(16^2) = 5/4;
Answer: k = 320; Intensity for 16 feet distance = 5/4